I need help with 6 problems, I have been trying to figure them out for 2 days now and have not been able too. Please help, here are the questions.
Find an equation of the line containing the given pair of points.
(1/4, -1/2) and (3/4, 6)
What is the equation of the line? y=
second question
Find an equation of the line having the given slope and containing the given point.
m=6/7, (1, -7)
The equation of the line is Y=
Third question
Find an equation of the line containing the given pair of points.
(-1, -2) and (-4, -6)
The equation of the line in slope-intercept from is y=
Fourth Question
Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b
(-4, 7); 4x=3y+2
The equation of the line is y=
Fifth question
Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b
(5, 7); 5x+y=4
the equation of the line is y=
Sixth question
Write an equation of the line containing the given point and perpendicular to the given line.
(5, -4); 4x+3y=8
The equation of the line is y=
I have no clue how to solve these problems, any help is greatly appreciated.
Thanks again

Answer by mananth(16949)

Find an equation of the line containing the given pair of points.
(1/4, -1/2) and (3/4, 6)
x1 y1 x2 y2
1/ 4 1/ 2 3/ 4 6

slope m = (y2-y1)/(x2-x1)
( 6 - 1/2 )/( 3/4 - 1/4 )
( 5.5 / 0.5 )
m= 11

Plug value of the slope and point ( 1/4 , 1/2 ) in
Y = m x + b
1/2 = 11/4 + b
b= 0.5 - 11/4
b= -9/ 4
So the equation will be
Y = 11 x -9/4
What is the equation of the line? y=
second question
Find an equation of the line having the given slope and containing the given point.
m=6/7, (1, -7)
m= 6/ 7

Plug value of the slope and point ( 1, -7
Y = m x + b
-7.00 = 6/7 + b
b= -7 - 6/7
b= -55/ 7
So the equation will be
Y = 6/7 x -55/7
The equation of the line is Y=
Third question
Find an equation of the line containing the given pair of points.
(-1, -2) and (-4, -6)
x1 y1 x2 y2
-1 -2 -4 -6

slope m = (y2-y1)/(x2-x1)
( -6 - -2 )/( -4 - -1
( -4 / -3 )
m= 4/ 3

Plug value of the slope and point ( -1 , -2 ) in
Y = m x + b
-2.00 = -4/3 + b
b= -2 - -4/3
b= - 2/ 3
So the equation will be
Y = 4/3 x - 2/3
The equation of the line in slope-intercept from is y=
Fourth Question
Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b
(-4, 7); 4x=3y+2
4 x + -3 y = 2
Find the slope of this line
make y the subject
-3 y = -4 x + 2
Divide by -3
y = 4/ 3 x - 2/3
Compare this equation with y=mx+b
slope m = 4/3
The slope of a line parallel to the above line will be the same
The slope of the required line will be 4/3
m= 4/3 ,point ( -4 , 7 )
Find b by plugging the values of m & the point in
y=mx+b
7 = -16/3 + b
b= 37/3
m= 4/3
Plug value of the slope and b in y = mx +b
The required equation is y = 4/3 x + 37/3


The equation of the line is y=
Fifth question
Write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b
(5, 7); 5x+y=4
5 x + 1 y = 4x+5y=2
Find the slope of this line
make y the subject
1 y = -5 x + 4
Divide by 1
y = -5 x + 4
Compare this equation with y=mx+b
slope m = -5
The slope of a line parallel to the above line will be the same
The slope of the required line will be -5
m= -5 ,point ( 5 , 7 )
Find b by plugging the values of m & the point in
y=mx+b
7 = -25.00 + b
b= 32
m= -5
Plug value of the slope and b in y = mx +b
The required equation is y = -5 x + 32
the equation of the line is y=
Sixth question
Write an equation of the line containing the given point and perpendicular to the given line.
(5, -4); 4x+3y=8
4 x + 3 y = 8
Find the slope of this line

3 y = -4 x + 8
Divide by 3
y = -1 1/ 3 x + 2.67
Compare this equation with y=mx+b
slope m = -4/ 3

The slope of a line perpendicular to the above line will be the negative reciprocal 3/4
m1*m2=-1


m= 3/4 ,point ( 5 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = 15/4 + b
b= -31/4
m= 3/4
The required equation is y = 3/ 4 x -31/ 4

The equation of the line is y=
m.ananth@hotmail.ca